OpenStudy (anonymous):

help @jim_thompson5910 For what x-value(s) does sin(x) = -1? (Points : 2) -270°, -90°, 90°, and 270° -270°, -90°, 0°, 90°, and 270° -90 and 270° -270° and 90°

jimthompson5910 (jim_thompson5910):

use the unit circle to find out which points on the unit circle have a y coordinate of 1, then record the angle for each point you found http://www.regentsprep.org/Regents/math/algtrig/ATT5/600px-Unit_circle_angles_svg.jpg

jimthompson5910 (jim_thompson5910):

y coordinate of -1 I meant to say

OpenStudy (anonymous):

how do you do it?

OpenStudy (anonymous):

@jim_thompson5910 i see 270

jimthompson5910 (jim_thompson5910):

good, so sin(270) = -1 this is the only angle in the interval from 0 degrees to 360 degrees where sine is -1

jimthompson5910 (jim_thompson5910):

all you need to do is list coterminal angles as well and you're done

OpenStudy (anonymous):

how do you do that?

jimthompson5910 (jim_thompson5910):

if you have any angle x, then a coterminal angle is x+360*n where n is any integer

jimthompson5910 (jim_thompson5910):

ex: if x = 30 and n = 1, then x+360n = 30+360*1 = 30+360 = 390 so the angle 390 degrees is coterminal to 30 degrees

OpenStudy (anonymous):

so its 270 and 90

jimthompson5910 (jim_thompson5910):

sin(90) = -1 ?? check the unit circle to confirm or deny this

OpenStudy (anonymous):

i see 60,45?

jimthompson5910 (jim_thompson5910):

locate the 90 degree angle what's the y coordinate of the point at this angle?

OpenStudy (anonymous):

its 1

jimthompson5910 (jim_thompson5910):

so sin(90) = 1 and NOT -1

jimthompson5910 (jim_thompson5910):

so if sin(x) = -1 then x cannot be 90 degrees

OpenStudy (anonymous):

-90,270 because sinx is negative in 3rd & 4th quadrants. sinx =-1=-sin 90=sin(-90),sin270 hence x is -90 & 270 in the given interval